OpenStudy (anonymous):

In the solution of the equation 5 - 3x = 2x + 9, 3x is added to the equation first. Which of the following should be done next? add -9 add -5 add -2x

OpenStudy (anonymous):

Can you write out the equation following the instruction " 3x is added to the equation first" ?

OpenStudy (anonymous):

What they really mean is 3x is added TO BOTH SIDES of the equation first.

OpenStudy (anonymous):

5-3x=2x+9 5=5x+9 -4=5x -4/5=x

OpenStudy (anonymous):

sodukogirl...hold off until we get a response pleae. and read the code of conduct for the website.

OpenStudy (anonymous):

3x=5 - 3x = 2x + 9+3x

OpenStudy (anonymous):

3x=

OpenStudy (anonymous):

*3x+

OpenStudy (anonymous):

5 - 3x = 2x + 9 + 3x +3x 5 = 5x + 9 :...................That's what we get after adding 3x to both sides

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

We trying to get the x alone on the right side. We need to get rid of that 9 and 5 on the right side so that x will be alone. Which one do I get rid of first? And how do I get rid of it?

OpenStudy (anonymous):

9 add a -9 to both sides

OpenStudy (anonymous):

Bingo! Good job.

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

how old r u